### v1

@@ 6.2 Proof of Theorem 2
A Convenient Bipartite Graph View of the Ownership Relation. To ease our discussion in the formal proofs regarding the optimal partition, we start by representing the overlapping ownership model as a bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$. The vertex sets $\mathcal{P},\mathcal{A},$ with $|\mathcal{P}|=n,|\mathcal{A}|=m$ correspond to the disjoint set of $n$ items (papers) and $m$ owners (authors), respectively. The edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{A}$ corresponds to the ownership relation: the owner $j$ owns the item $i$, if and only if there is an edge $(p_{i},a_{j})\in\mathcal{E}$ between the vertices $p_{i}\in\mathcal{P}$ and $a_{j}\in\mathcal{A}$. The number of edges $|E|=\sum_{j\in[m]}|\mathcal{I}^{j}|$ is precisely the $N$ in the theorem statement, which captures the order of the problem’s input size. Let $E=(e_{i}^{j})_{m\times n}$ be the $0-1$ bi-adjacency matrix of the ownership in which $e_{i}^{j}=1$ for each edge $(p_{i},a_{j})\in\mathcal{E}$ — as is illustrated in Table 2 and 3. Hence, the set of items owned by $j$ is exactly the neighbor set of vertex $a_{j}$, i.e., $\mathcal{I}^{j}=\{p_{i}\in\mathcal{P}|e_{i}^{j}=1\}$.
@@
We now formalize the partition-based scheme in Mechanism 2 using the bipartite graph characterization. Specifically, we consider the partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ on the vertex set $\mathcal{P}$ of the bipartite graph. By definition of a partition, we have $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=\mathcal{P}$ and ${\mathcal{S}}_{k^{\prime}}\cap{\mathcal{S}}_{k}=\varnothing,\forall k\neq k^{\prime}\in[K]$. For each partition block ${\mathcal{S}}_{k}$, we denote ${\mathcal{T}}_{k}=\{j\in\mathcal{A}|{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$ as the set of owners who owns all items in ${\mathcal{S}}_{k}$. For the blocks with $|{\mathcal{S}}_{k}|\leq 1$ or $|{\mathcal{T}}_{k}|\leq 0$, the mechanism simply uses the raw review score. Otherwise, by construction, the sub-bipartite-graph of $\mathcal{G}$ induced by each $({\mathcal{S}}_{k},{\mathcal{T}}_{k})$ and their incidence edges $\{(p_{i},a_{j})\in\mathcal{E}|p_{i}\in{\mathcal{S}}_{k},a_{j}\in{\mathcal{T}}_{k}\}$ forms a complete bipartite graph, i.e., a complete overlapping ownership. Therefore, we can also visualize the partition of the partially overlapping ownership using the bi-adjacency matrix illustrated in Table 3.
@@
Proof of Theorem 2.
Pick any ownership instance represented as a bipartite graph $\mathcal{G}$ and any weight function $w\in\mathcal{W}$. Let $\operatorname{OPT}(\mathcal{G}):=\operatorname{obj}({\mathcal{S}}^{*})$, where ${\mathcal{S}}^{*}=\{{\mathcal{S}}_{k}^{*}\}_{k=1}^{K}$ is the optimal $1$-strong partition of $\mathcal{G}$. Consider the greedy algorithm that keeps picking the largest residual item set owned by some author, until all items are assigned to some block, as described in Algorithm 3. Let ${\mathcal{S}}$ denote the partition obtained by Algorithm 3. By construction, ${\mathcal{S}}$ is 1-strong — i.e., there is at least one owner who owns all items of each item subset in ${\mathcal{S}}$. We let $\operatorname{ALG}(\mathcal{G}):=\operatorname{obj}({\mathcal{S}})$ be the objective value of the partition obtained by Algorithm 3. The remainder of the proof will argue that the output ${\mathcal{S}}$ simultaneously satisfies,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=\frac{\operatorname{obj}({\mathcal{S}})}{\operatorname{obj}({\mathcal{S}}^{*})}\geq\inf\big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\big{\}},$
@@
Let $j$ be the very first owner picked by Algorithm 3 on the bipartite graph instance $\mathcal{G}$ with $|\mathcal{A}|=k+1$. Thus, the first block (i.e., item set) picked by the algorithm must be $\mathcal{I}^{j}$. Let ${\mathcal{S}}^{*}$ be the optimal partition of $\mathcal{G}$ under weight $w$. Note that the algorithm output ${\mathcal{S}}$ is independent of $w$, but ${\mathcal{S}}^{*}$ typically depends on $w$. For each block ${\mathcal{S}}_{k}^{*}\in{\mathcal{S}}^{*}$, let $p_{k}^{*}:=\left|{\mathcal{S}}_{k}^{*}\right|$ be the number of items in this block and $a_{k}:=\left|{\mathcal{S}}^{*}_{k}\cap\mathcal{I}^{j}\right|$ be the number of items covered by the greedy algorithm’s first choice $\mathcal{I}^{j}$. Hence, $\left|\mathcal{I}^{j}\right|=\sum_{k=1}^{K}a_{k}$. We can then bound the approximation ratio of the greedy algorithm by decomposing the objective value as follows,
@@
We will bound the approximation ratio by considering the two parts $\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$ and $\frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}$ separately, and then apply the mediant inequality to derive the lower bound. The second part can be easily bounded according to the induction hypothesis. However, the first part could be unbounded in general (this may happen when $p^{*}_{k}$’s are all at most 1). Fortunately, the analysis for such special case can be carried out separately. Specifically, if $\mathop{max}_{k\in[K]}p^{*}_{k}\leq 1$ in $\mathcal{G}$, by convexity of $w$ function, this implies that any 1-strong partition of the items are at least as good as ${\mathcal{S}}^{*}$ and thus the greedy algorithm is optimal.
@@
Finally, for running time analysis, Algorithm 3 can be efficiently implemented in almost linear time using appropriately chosen data structures. Specifically, we employ the red-black tree data structure which can retrieve the maximum element $j^{*}=\arg\mathop{max}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$, i.e., the maximal residual paper sets among all authors, using $O(\log m)$ time. We will then update the paper sets of all the authors who have at least one paper in $\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$. The number of operations is at most the number of edges connected to ${\mathcal{S}}^{j^{*}}$ in the bipartite graph. We then update the red-black tree structure using the new sizes of the residual paper sets of the involved papers. Note that the update of each author’s paper set size takes $O(\log m)$ time by first deleting it in the red-black tree and then re-inserting it. The above recursion takes at most $m$ rounds since each round at least one author’s residual paper set will become empty and there are only $m$ authors. In total, the total running time is thus $O(|\mathcal{E}|\log m)$.
@@
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}$
Proof of Lemma 4.
Any $w\in\mathcal{W}$ is convex, so its left derivative $w^{\prime}_{-}$ is non-deceasing such that $w(p^{*}_{k})-w(p^{*}_{k}-a_{k})\leq w^{\prime}_{-}(p_{k}^{*})a_{k}$. In addition, $w^{\prime}_{-}(p^{*}_{k})\geq 0$ since $w(0)=0,w(p^{*}_{k})\geq 0$. Let $\sum_{k=1}^{K}a_{k}=a$, we have
@@
Since $a\geq\mathop{max}_{k\in[K]}p^{*}_{k}=p$ and thus $w^{\prime}_{-}(p_{k}^{*})\leq w^{\prime}_{-}(a)$ for any $k$ by convexity of $w$. Since $w(p)>0$ and $w(0)=0$, we must have $0<w^{\prime}_{-}(p)\leq w^{\prime}_{-}(x)$ for any $x\geq p$. Therefore,
@@
6.2.1 Tightness of Greedy’s Approximation Ratio in the Monomial Class
@@
every picture/.style=line width=0.75pt
[x=0.75pt,y=0.75pt,yscale=-1,xscale=1]
[…]
@@
Meanwhile, despite the hardness result for the size-based objective in Proposition 2, we would like to point out that the efficient algorithm for optimal partition under the $\alpha$-th degree polynomial objective may exist. In particular, consider the following instance with $m=n+1$, where the optimal partition cannot realize the set-cover problem for the hardness reduction. The owner $i$ owns the items $\{2i-1,2i\}$,$\forall i\in[n]$ and owner $n+1$ owns the items $\{1,3,\cdots,2n-1\}$. In this case, the optimal partition is not the minimum set cover, since
@@
7 Discussion
In this work, we have established a generalized Isotonic Mechanism designed to enhance peer review in machine learning conferences, specifically addressing the practical context where multiple authors contribute to various papers. We demonstrate the truthfulness of our mechanism and design a simple polynomial-time algorithm to partition the ownership sets, which achieves a constant ratio approximation of the optimal partition that is computationally hard to find. This facilitates the practical implementation of our mechanism.
Nonetheless, our mechanism is not without limitations, which present opportunities for future investigations. A natural extension lies in adapting the mechanism to accommodate review scores drawn from exponential family distributions [26]. Additionally, the introduction of randomness into authors’ private information presents an intriguing research direction. This modeling perspective could more accurately reflect the reality that authors may assign different values to their co-authored submissions. Lastly, recognizing the potential heterogeneity of the utility function poses a significant, yet crucial challenge. In this context, an author’s value of a paper could depend on factors such as authorship order, the length of the author’s list, and her relative contributions to the paper [7].
Acknowledgments
This work was supported in part by NSF through CCF-2132506 and CCF1934876, ARO through W911NF-23-1-0030, Analytics at Wharton, and Wharton AI and Analytics for Business.
Appendix A Appendix
Proof of Proposition 1.
Pick any bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ from the ownership instance $\mathcal{O}=\{\mathcal{I}^{j}\}_{j\in[m]}$. Suppose we were to find $L$-strong partition ${\mathcal{S}}$ in $\mathcal{G}$. We can construct a bipartite graph $\mathcal{G}^{\prime}=(\mathcal{A}^{\prime},\mathcal{P}^{\prime},\mathcal{E}^{\prime})$, where the item set remains the same $\mathcal{P}^{\prime}=\mathcal{P}$, the owner set $\mathcal{A}^{\prime}$ corresponds to all the $L$ element subsets of $\mathcal{A}$, $\mathcal{A}^{\prime}=\{\ell|u^{\ell}\subseteq\mathcal{A},|u^{\ell}|=L\}$ and the edge set captures the commons items of every $L$ owners, =$\mathcal{E}^{\prime}=\{(i,\ell)|i\in\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}\}$. We show that for any $L$-strong partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},{\mathcal{S}}_{2},\dots,{\mathcal{S}}_{K}\}$ of $\mathcal{G}$, there is an equivalent $1$-strong greedy partition ${\mathcal{S}}^{\prime}=\{{\mathcal{S}}^{\prime}_{1},{\mathcal{S}}^{\prime}_{2},\dots,{\mathcal{S}}^{\prime}_{K}\}$ of $\mathcal{G}^{\prime}$ such that ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in K$.
The proof is a straightforward verification of this bipartite graph construction. Pick any ${\mathcal{S}}_{k}\in{\mathcal{S}}$ with $|{\mathcal{T}}_{k}|\geq L$ in $\mathcal{G}$, there exists an ${\mathcal{S}}^{\prime}_{k}$ with $|{\mathcal{T}}^{\prime}_{k}|=1$ in $\mathcal{G}^{\prime}$. That is, we have ${\mathcal{T}}^{\prime}_{k}=\{\ell\}$ for some $u^{\ell}\subseteq{\mathcal{T}}_{k}$, then ${\mathcal{S}}^{\prime}_{k}={\mathcal{S}}_{k}\subseteq\bigcap_{j\in u^{\ell}}\mathcal{I}^{j}$ by construction. Therefore, ${\mathcal{S}}_{k}={\mathcal{S}}^{\prime}_{k},\forall k\in[K]$. ∎
Proof of Proposition 2.
We consider the weight function $w(x)=\mathop{max}\{x,1\}-1$. For any partition with $K$ blocks, the objective function is $\operatorname{obj}({\mathcal{S}})=n-K$. We show that finding the optimal partition under such an objective function is at least as hard as finding the minimum set. We prove via a standard hardness reduction argument: for any set cover instance, we can construct a bipartite graph instance in which the partition that maximizes the objective function is $\operatorname{obj}({\mathcal{S}})=n-K$ can be turned into a minimum set cover.
Pick arbitrary set cover instance with $m$ subsets of $n$ item, $\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{m}$. We construct a bipartite graph $\mathcal{G}=(\mathcal{A},\mathcal{P},\mathcal{E})$ where $\mathcal{A}=[m],\mathcal{P}=[n]$ and =$\mathcal{E}=\{(p_{i},a_{j})|p_{i}\in\mathcal{V}_{j},j\in[m]\}$. There exists an optimal partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$, if and only if the minimum set cover has size $K^{*}$. For the “if” direction, we construct a partition ${\mathcal{S}}^{*}$ with $\operatorname{obj}({\mathcal{S}}^{*})=n-K^{*}$ from the set cover of size $K^{*}$. For each subset selected by the set cover, we construct a partition block with items in the subset minus the items already in the partition. This partition must be valid and have $K^{*}$ blocks. For the “only if” direction, we construct the set cover of size $K^{*}$ from the optimal partition ${\mathcal{S}}^{*}$. For every block ${\mathcal{S}}_{k}^{*}$ of the optimal partition ${\mathcal{S}}^{*}$, we pick any vertex $a_{j}\in{\mathcal{T}}_{k}^{*}$ from the corresponding owner set. Based on the $*K*$ distinct vertices, we select the subsets with the same indices, and it forms a set cover of all $n$ items by construction. ∎

### v3

@@ Appendix E Proof of Theorem 3
A Convenient Bipartite Graph View of the Ownership Relation. To ease our discussion in the formal proofs regarding the optimal partition, we start by representing the overlapping ownership model as a bipartite graph $\mathcal{G}=(\mathcal{P},\mathcal{A},\mathcal{E})$. The vertex sets $\mathcal{P},\mathcal{A},$ with $|\mathcal{P}|=n,|\mathcal{A}|=m$ correspond to the disjoint set of $n$ items (papers) and $m$ owners (authors), respectively. The edge set $\mathcal{E}\subseteq\mathcal{P}\times\mathcal{A}$ corresponds to the ownership relation: the owner $j$ owns the item $i$, if and only if there is an edge $(p_{i},a_{j})\in\mathcal{E}$ between the vertices $p_{i}\in\mathcal{P}$ and $a_{j}\in\mathcal{A}$. The number of edges $|E|=\sum_{j\in[m]}|\mathcal{I}^{j}|$ is precisely the $N$ in the theorem statement, which captures the order of the problem’s input size. Let $E=(e_{i}^{j})_{m\times n}$ be the $0-1$ bi-adjacency matrix of the ownership in which $e_{i}^{j}=1$ for each edge $(p_{i},a_{j})\in\mathcal{E}$. Hence, the set of items owned by $j$ is exactly the neighbor set of vertex $a_{j}$, i.e., $\mathcal{I}^{j}=\{p_{i}\in\mathcal{P}|e_{i}^{j}=1\}$.
@@
We now formalize the partition-based scheme in Mechanism 2 using the bipartite graph characterization. Specifically, we consider the partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ on the vertex set $\mathcal{P}$ of the bipartite graph. By definition of a partition, we have $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=\mathcal{P}$ and ${\mathcal{S}}_{k^{\prime}}\cap{\mathcal{S}}_{k}=\varnothing,\forall k\neq k^{\prime}\in[K]$. For each partition block ${\mathcal{S}}_{k}$, we denote ${\mathcal{T}}_{k}=\{j\in\mathcal{A}|{\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}\}$ as the set of owners who owns all items in ${\mathcal{S}}_{k}$. For the blocks with $|{\mathcal{S}}_{k}|\leq 1$ or $|{\mathcal{T}}_{k}|\leq 0$, the mechanism simply uses the raw review score. Otherwise, by construction, the sub-bipartite-graph of $\mathcal{G}$ induced by each $({\mathcal{S}}_{k},{\mathcal{T}}_{k})$ and their incidence edges $\{(p_{i},a_{j})\in\mathcal{E}|p_{i}\in{\mathcal{S}}_{k},a_{j}\in{\mathcal{T}}_{k}\}$ form a complete bipartite graph, i.e., a complete overlapping ownership. Therefore, we can also visualize the partition of the partially overlapping ownership using the bi-adjacency matrix.
@@
Proof.
Proof of Theorem 3 Pick any ownership instance represented as a bipartite graph $\mathcal{G}$ and any wellness function $w\in\mathcal{W}$. Let $\operatorname{OPT}(\mathcal{G}):=\operatorname{obj}(\mathfrak{S}^{*})$, where $\mathfrak{S}^{*}=\{{\mathcal{S}}_{k}^{*}\}_{k=1}^{K}$ is the optimal $1$-strong partition of $\mathcal{G}$. Consider the greedy algorithm that keeps picking the largest residual item set owned by some owner, until all items are assigned to some block, as described in Algorithm 4. Let $\mathfrak{S}$ denote the partition obtained by Algorithm 4. By construction, $\mathfrak{S}$ is 1-strong — i.e., there is at least one owner who owns all items of each block in $\mathfrak{S}$. We let $\operatorname{ALG}(\mathcal{G}):=\operatorname{obj}(\mathfrak{S})$ be the objective value of the partition obtained by Algorithm 4. The remainder of the proof will argue that the output $\mathfrak{S}$ simultaneously satisfies,
$\frac{\operatorname{ALG}(\mathcal{G})}{\operatorname{OPT}(\mathcal{G})}=\frac{\operatorname{obj}(\mathfrak{S})}{\operatorname{obj}(\mathfrak{S}^{*})}\geq\inf\big{\{}\frac{w(x)}{w^{\prime}_{-}(x)x}\big{|}w^{\prime}_{-}(x)>0,x\geq 2\big{\}},$
@@
Let $j$ be the very first owner picked by Algorithm 4 on the bipartite graph instance $\mathcal{G}$ with $|\mathcal{A}|=k+1$. Thus, the first block (i.e., item set) picked by the algorithm must be $\mathcal{I}^{j}$. Let $\mathfrak{S}^{*}$ be the optimal partition of $\mathcal{G}$ under weight $w$. Note that the algorithm output $\mathfrak{S}$ is independent of $w$, but $\mathfrak{S}^{*}$ typically depends on $w$. For each block ${\mathcal{S}}_{k}^{*}\in\mathfrak{S}^{*}$, let $p_{k}^{*}:=\left|{\mathcal{S}}_{k}^{*}\right|$ be the number of items in this block and $a_{k}:=\left|{\mathcal{S}}^{*}_{k}\cap\mathcal{I}^{j}\right|$ be the number of items covered by the greedy algorithm’s first choice $\mathcal{I}^{j}$. Hence, $\left|\mathcal{I}^{j}\right|=\sum_{k=1}^{K}a_{k}$. We can then bound the approximation ratio of the greedy algorithm by decomposing the objective value as follows,
@@
We will bound the approximation ratio by considering the two parts $\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}$ and $\frac{\operatorname{ALG}(\mathcal{G}\setminus j)}{\operatorname{OPT}(\mathcal{G}\setminus j)}$ separately, and then apply the mediant inequality to derive the lower bound. The second part can be easily bounded according to the induction hypothesis. However, the first part could be unbounded in general (this may happen when $p^{*}_{k}$’s are all at most 1). Fortunately, the analysis for such special case can be carried out separately. Specifically, if $\mathop{max}_{k\in[K]}p^{*}_{k}\leq 1$ in $\mathcal{G}$, by convexity of $w$ function, this implies that any 1-strong partition of the items are at least as good as $\mathfrak{S}^{*}$, thus the greedy algorithm is optimal.
@@
Finally, for running time analysis, Algorithm 4 can be efficiently implemented in almost linear time using appropriately chosen data structures. Specifically, we employ the red-black tree data structure which can retrieve the maximum element $j^{*}=\arg\mathop{max}_{j\in[m]}|\mathcal{I}^{j}\setminus\overline{\mathcal{I}}|$, i.e., the maximal residual item sets among all owners, using $O(\log m)$ time. We will then update the item sets of all the owners who have at least one item in $\mathcal{I}^{j^{*}}\setminus\overline{\mathcal{I}}$. The number of operations is at most the number of edges connected to removed items from $j^{*}$ in the bipartite graph. We then update the red-black tree structure using the new sizes of the residual item sets of the involved items. Note that the update of each owner’s item set size takes $O(\log m)$ time by first deleting it in the red-black tree and then re-inserting it. The above recursion takes at most $m$ rounds since each round at least one owner’s residual item set will become empty and there are only $m$ owners. In total, the total running time is thus $O(|\mathcal{E}|\log m)$.
@@
$\frac{w\left(\sum_{k=1}^{K}a_{k}\right)}{\sum_{k=1}^{K}w(p^{*}_{k})-\sum_{k=1}^{K}w(p^{*}_{k}-a_{k})}\geq\inf_{x\geq p}\frac{w(x)}{w^{\prime}_{-}(x)x}.$
Proof.
Proof of Lemma 6 Any $w\in\mathcal{W}$ is convex, so its left derivative $w^{\prime}_{-}$ is non-deceasing such that $w(p^{*}_{k})-w(p^{*}_{k}-a_{k})\leq w^{\prime}_{-}(p_{k}^{*})a_{k}$. In addition, $w^{\prime}_{-}(p^{*}_{k})\geq 0$ since $w(0)=0,w(p^{*}_{k})\geq 0$. Let $\sum_{k=1}^{K}a_{k}=a$, we have
@@
Since $a\geq\mathop{max}_{k\in[K]}p^{*}_{k}=p$, thus $w^{\prime}_{-}(p_{k}^{*})\leq w^{\prime}_{-}(a)$ for any $k$ by convexity of $w$. Since $w(p)>0$ and $w(0)=0$, we must have $0<w^{\prime}_{-}(p)\leq w^{\prime}_{-}(x)$ for any $x\geq p$. Therefore,
@@
E.1 Tightness of Greedy’s Approximation Ratio in the Monomial Class
@@
Meanwhile, despite the hardness result for the size-focused objective in Proposition 2, we would like to point out that the efficient algorithm for optimal partition under the $\alpha$-th degree polynomial objective may exist. In particular, consider the following instance with $m=n+1$, where the optimal partition cannot realize the set-cover problem for the hardness reduction. The owner $i$ owns the items $\{2i-1,2i\}$,$\forall i\in[n]$ and owner $n+1$ owns the items $\{1,3,\cdots,2n-1\}$. In this case, the optimal partition is not the minimum set cover, since
@@
Appendix F Additional Experiments
In Figure 9, we show how MSE improvement influences the accuracy of more stringent paper selections, e.g., spotlight (top 5%) and oral (top 1.5%) presentations. We can observe that the calibration performance has similar trend of improvement as the poster acceptance.
(a) ICLR 2021 Oral
(b) ICLR 2022 Oral
(c) ICLR 2023 Oral
(d) ICLR 2021 Spotlight
(e) ICLR 2022 Spotlight
(f) ICLR 2023 Spotlight
Figure 9: The percentile precision of Isotonic Mechanisms for Oral and Spotlight under varying level review noise $\sigma$ in ICLR.